# Phase transitions and spontaneous symmetry breaking in the renormalized Ginzburg-Landau theory

### v1

@@ II.2 Derivation of new quadratic coefficients
For $d=2$, the expression for the renormalized quadratic coefficient is given by $r(d=2,T)=\eta\left(\frac{T}{T_{c}}\right)+r_{0}(T-T_{c})$. This expression is relatively straightforward, similar to the thermodynamic quantities derived from it. The temperature at which $r(d=2,T)$ changes sign decreases as $r(d=2,T)$ increases and is significantly influenced by the nature of fluctuations in the system (see Eq. (12) for $d=2$)). For large values of $\eta$, $T_{c_{2D}}$ can approach zero, thereby confirming the Mermin-Wagner theorem. For $d>2$, the solutions are nearly the same as those obtained from standard theory, with the main difference being that thermodynamic parameters depend on the intrinsic nature of the systems involved. While Landau’s theory does not allow us to predict the exact transition temperature, it does facilitate an analysis of the system’s behavior near the transition. In our model, we emphasize that the mean-field transition temperature represents a characteristic thermal scale rather than the true transition temperature for systems with dimensions $d\leq 2$. This highlights the necessity of modifying Landau’s theory to account for the crucial role of dimensionality in the manifestation and analysis of transitions, which is important for understanding concrete phenomena. Several open questions in phase transitions and critical phenomena remain pertinent. A key question is whether all low-temperature phase effects arise solely from fluctuation-induced order parameters. Given that pre-transitional behaviors provide insights into the physical mechanisms underlying phase transitions, our approach diverges from the standard one and necessitates numerical investigation of Eqs. (8) and (9) to better understand the contribution of the correction term $\Omega(d,T)$.
III Renormalized thermodynamic quantities
In the continuum limit, the Ginzburg-Landau theory provides a generalized partition function $Z$, which is derived by integrating over all possible configurations $\mathcal{D}\phi(\textbf{r})$ of the order parameter [26].
@@ III Renormalized thermodynamic quantities
The correlation function relating the order parameter and density at different positions is given by
@@
For more information on calculating thermodynamic functions, we refer readers to Ref. [4]. We decompose $\phi(\textbf{r})$ into its Fourier components $\phi(\textbf{q})$ as follows:
$\phi(\textbf{q})=\frac{1}{V}\int d^{d}\textbf{r}\phi(\textbf{r})\exp\{-i\textbf{q}\cdot\textbf{r}\}.$ (31)
By applying Fourier modes from Equation (17) in the Ginzburg-Landau functional without an external field, we can establish a proportional relationship between the variance of the order parameter (which represents fluctuations) and the average change of that order parameter when a parameter changes. This relationship is known as the fluctuation-dissipation theorem, and it can be expressed as follows:
In the absence of an external field ($h=0$) and referring to Eq. (5), one can obtain the susceptibility modified by fluctuations in the following manner:
$\chi^{-1}\Big{|}_{h\rightarrow 0}=\left\{\begin{array}[]{ll}\vspace{0,5cm}r_{0}(T-T_{c})+\Omega(d,T),&\textrm{for}\phantom{.}T>T_{c_{d}}\\ 2\Big{(}r_{0}(T_{c}-T)-\Omega(d,T)\Big{)}.&\textrm{for}\phantom{.}T<T_{c_{d}}.\\ \end{array}\right.$ (35)
the function $\chi^{-1}(T)$ depend on the dimensional and intrinsic parameters of the system. There is a challenge when accurately evaluating certain thermodynamic quantities, such as specific heat, within the Ginzburg-Landau model. In some cases, it is necessary to apply the Gaussian approximation above the transition temperature. This approach is achieved by omitting the $|\phi|^{4}$-term, which is commonly linked to the neglect of scattering state contributions in Equation (6). As a result, when integrating over all fluctuation degrees of freedom, the renormalized specific heat, without considering an external field, can be factorized as follows:
$C_{p}=\left\{\begin{array}[]{ll}\vspace{0,5cm}\frac{1}{2}\gamma_{d}{(\xi^{+}_{0})^{-d}}\Big{[}\frac{T}{T_{c}}\Big{]}^{2}\Big{(}1+T_{c}\frac{\partial\kappa(d,T)}{\partial T}\Big{)}^{2}\tau_{d}^{-\alpha},&\textrm{for}\phantom{.}T<T_{c_{d}}\\ \Delta{C_{0}}(d,T_{c_{d}})+\gamma_{d}{(\xi^{-}_{0})^{-d}}\Big{[}\frac{T}{T_{c}}\Big{]}^{2}\Big{(}1+T_{c}\frac{\partial\kappa(d,T)}{\partial T}\Big{)}^{2}\Big{|}\tau_{d}\Big{|}^{-\alpha^{\prime}},&\textrm{for}\phantom{.}T<T_{c_{d}},\\ \end{array}\right.$ (38)
$\kappa(d,T)=\frac{\Omega(d,T)}{r_{0}T_{c}},$ (39)
IV Results and discussions
$\Delta{C_{0}}(d,T)=\frac{r^{2}_{0}T}{4b_{0}}\Bigg{[}\bigg{(}1+T_{c}\frac{\partial{\kappa(d,T)}}{\partial{T}}\bigg{)}^{2}+T^{2}_{c}\Big{(}\epsilon+\kappa(d,T)\Big{)}\frac{\partial^{2}{\kappa(d,T)}}{{\partial{T}}^{2}}\Bigg{]}.$ (40)
In the expression of Eq. (22), when $\Big{(}\epsilon+\kappa(d,T)\Big{)}\Big{|}_{T=T_{c_{d}}}=0$ vanishes at the renormalized transition point, the quantity $\bigg{(}1+T_{c}\frac{\partial{\kappa(d,T)}}{\partial{T}}\Big{|}_{T=T_{c_{d}}}\bigg{)}^{2}$ provides information about the magnitude of the jump in the specific heat at that same point. If there are significant contributions from scattering states, the height of this jump can either increase or decrease substantially. Additionally, the interaction between $r_{0}(T-T_{c})$ and $\Omega(d,T)$ may result in inhomogeneities in the sample, which can potentially lead to the disappearance of the specific heat jump altogether. Hence, the anomalous part of the specific heat at the renormalized critical point is defined by this relationship.
By setting $\varepsilon=\frac{\eta}{r_{0}T_{c}}$, we can easily establishe the following result
$\Delta{C_{0}}(d,T_{c_{d}})=\left\{\begin{array}[]{lll}0,&\textrm{for}\phantom{...}d=1\\ (1+\varepsilon)\Delta{C_{0}},&\textrm{for}\phantom{...}d=2\\ 0\phantom{..}\textrm{or}\phantom{..}4(1+\varepsilon^{2})^{2}\Delta{C_{0}},&\textrm{for}\phantom{...}d=3.\\ \end{array}\right.$ (45)
where $\Delta{C_{0}}=\frac{r^{2}_{0}T_{c}}{4b_{0}}$ is the mean-field specific heat jump. The renormalized specific heat jump is influenced by the system’s dimension and intrinsic parameters. Self-consistent calculations of Eq. (9) yield corrections that describe either a disappearance, linear growth, or quasi-exponential growth of the specific heat jump at the renormalized critical point. Linear growth is observed in two-dimensional systems, while exponential growth of the specific heat jump is seen in three-dimensional systems. Notably, one-dimensional systems do not exhibit a specific heat jump, and three-dimensional systems may also lack this jump under certain conditions. In fact, by solving Equation (9) for $d=3$, we find two solutions; one of these leads to the absence of the specific heat jump at the critical temperature. This phenomenon occurs due to the competition between $r_{0}(T-T_{c})$ and $\Omega(d,T)$, which results in competing phases within the system and, consequently, a competition between the resulting order parameters. Additionally, the disappearance of the specific heat jump in superconducting systems can be attributed to a crossover in the type of order parameters, as discussed in Refs. [10], which leads to a significant loss of entropy at the critical temperature [9, 27]. Exponential growth of the specific heat jump has been observed in families of $\beta$-pyrochlore oxide superconductors [28, 29, 30], which can be explained by a significant lattice contribution coupled with strong fluctuations in the order parameter. However, certain families of high-$T_{c}$ superconductors, such as $TlBaCaCuO$ and $BiSrCaCuO$, do not exhibit a conventional specific heat jump at the critical temperature ($T_{c}$) [11]. This lack of a clear jump can be attributed to the presence of a partial gap in the electronic density of states at the Fermi level [31]. Another potential explanation involves the smearing of $T_{c}$ due to sample inhomogeneities and the competition or interplay between superconductivity and other electronic phases, such as antiferromagnetism. Qualitatively, the results from Equation (24) suggest that a small value of $\eta$ can either cancel out or significantly enhance the jump in specific heat at the critical temperature.
V Conclusion
Acknowledgments

### v2

@@ II.2 Dimension-dependent renormalized quadratic coefficient and treatment
where $\epsilon=\frac{\eta}{r_{0}T_{c}}$ defines the scaled fluctuation correction ratio term. For $d$ = 2, the expression for the renormalized quadratic coefficient is given by $r(d=2,T)=\eta\left(\frac{T}{T_{c}}\right)+r_{0}(T-T_{c})$. This expression is relatively straightforward, similar to the thermodynamic quantities derived from it. The temperature at which $r(d=2,T)$ changes sign decreases as $r(d=2,T)$ increases and is significantly influenced by the nature of fluctuations in the system (see Eq. (27) for $d$ = 2)). For large values of $\eta$, $T_{c_{2D}}$ can approach zero, thereby confirming the Mermin-Wagner theorem [41]. In the following paragraph, we will discuss in detail the behavior of the transition temperature in two-dimensional systems. For $d>2$, the solutions are nearly the same as those obtained from standard theory, with the main difference being that thermodynamic parameters depend on the intrinsic nature of the systems involved. While Landau’s theory does not allow us to predict the exact transition temperature, it does facilitate an analysis of the system’s behavior near the transition. In our model, we emphasize that the mean-field transition temperature represents a characteristic thermal scale rather than the true transition temperature for systems with dimensions $d\leq 2$. This highlights the necessity of modifying Landau’s theory to account for the crucial role of dimensionality in the manifestation and analysis of transitions, which is important for understanding concrete phenomena. Several open questions in phase transitions and critical phenomena remain pertinent. A key question is whether all low-temperature phase effects arise solely from fluctuation-induced order parameters. Given that pre-transitional behaviors provide insights into the physical mechanisms underlying phase transitions, our approach diverges from the standard one and necessitates numerical investigation of Eqs. (14) and (16) to better understand the contribution of the correction term $\Omega(d,T)$.
III General framework and Renormalized thermodynamic quantities
Having determine the fluctuations contribution at quadratic order to the free energy, it is necessary to evaluate the fluctuation contribution to some thermodynamic quantities. As well-known, the universality that characterizes the behavior of thermodynamic functions close to a second order phase transition appears as one of the fascinating aspects of critical phenomena in statistical mechanics [5, 19, 37]. By applying Fourier modes as presented in the appendix in the Ginzburg-Landau functional and in the absence of an external field, we can establish a proportional relationship between the variance of the order parameter (which represents fluctuations) and the average change of that order parameter when the external field changes. This relationship, known as the fluctuation-dissipation theorem, can be expressed as follows:
which relates the susceptibility $\chi$ and the correlation function $C$. More generally, this fluctuation-response relation is valid for weak fluctuations since linear response is assumed. Importantly, in the absence of an external field ($h=0$) and referring to Eq. (11), one can obtain the susceptibility modified by fluctuations in the following manner:
[…]
Importantly, the temperature -dependent inverse susceptibility function $\chi^{-1}(T)$ depend explicitly on the dimensional and intrinsic parameters of the system. There is a challenge when accurately evaluating certain thermodynamic quantities, such as specific heat, within the Ginzburg-Landau model. In some cases, it is necessary to apply the Gaussian approximation above the transition temperature. This approach is achieved by omitting the $|\phi|^{4}$-term, which is commonly linked to the neglect of scattering state contributions in Eq. (12). As a result, when integrating over all fluctuation degrees of freedom, the renormalized specific heat, without considering an external field, can be factorized as follows:
[…]
$\kappa(d,T)=\frac{\Omega(d,T)}{r_{0}T_{c}},$ (35)
Indeed, at this extend following the more general result known as the Mermin-Wagner theorem [41], which states that there is no spontaneous breaking of a continuous symmetry in systems with short-range interactions in dimensions $d\leq 2$, as the phase fluctuations become asymptotically large, leading to the corollary existence of the borderline dimensionality of two, known as the lower critical dimension $d_{l}$. The renormalized approach effectively illustrates and reaffirms the Mermin-Wagner theorem [41], particularly in one-dimensional systems. In two-dimensional systems, while recovering significant critical fluctuations, one observes a variation of the true critical point $T_{c_{2D}}$ between 0 K and $T_{c}$, i.e., $0\leq T_{c_{2D}}\leq T_{c}$.
Figure 4: Profile of fluctuation-correction dependence of the two-dimensional critical temperature according to Eq. (27) for $T_{c}$ = 2.5 K. Here, the scaled fluctuation correction ratio $\epsilon$ defines the relative temperature distance from $T_{c}$.
Fig. 4 shows the temperature behavior of the true critical temperature according to the scaled fluctuation-correction term $\epsilon$. Here, a careful treatment bring a notable outcome with respect to the Mermin-Wagner theorem, which asserts that a continuous symmetry cannot be broken in two dimensions. These findings show that the presence of fluctuations leads to a decrease in the degree of order in two-dimension (dd= 2) and completely destroys it in dimensions less than two ($d<2$). The temperature behavior illustrated in Fig. 4 suggests a transition from classical two-dimensional fluctuations at $\epsilon=0$ to quantum two-dimensional critical fluctuations that occur over a broad temperature range when $\epsilon\gg 1$. For each value of the $\epsilon$ term, we can accurately determine the fluctuation term across a wide range of temperatures. Quantum fluctuations play a crucial role in driving quantum phase transitions, such as the quantum melting of Wigner crystals into Fermi liquids in electron systems. However, the effects of these fluctuations on superconducting systems close to absolute zero, especially regarding the superconductor-insulator/metal transition, remain an open question. Moreover, the parameters associated with this approach depend significantly on both the dimension and the size of the sample, but in ways that differ from those derived from the mean-field approach. As a result, the renormalized approach appears to be well-suited for model calculations in various critical systems [25, 27, 2, 42].
From three-dimensional systems, standard and renormalization models are in good agreement, especially regarding the position of the critical point, with, however, critical behaviors being qualitatively and quantitatively different, as we will see in the following paragraph, concerning the specific heat jump $\Delta{C_{p}}$. However, it appears that Ginzburg-Landau models which belong to the universality class exhibit an upper critical dimension $d_{u}$ of four, through the so-called the Ginzburg Criterion which establishes the importance limit of fluctuations and which has to be treated carefully. The impact of new quadratic coefficients in three dimensional systems is analyzed in greater detail elsewhere, specifically regarding the non-monotonic variation of the specific heat jump, $\Delta{C_{p}}/T_{c}$, across the transition for various high-temperature superconductors [43]. For instance, in YBCO synthesized as $YBa_{2}Cu_{3}O_{7-\delta}$ with $0\leq\delta\leq 0.18$, variations in oxygen stoichiometry can affect the density of states at the Fermi level and, consequently, the magnitude of $\Delta{C_{p}}$. Observations of transitions from relatively weak three-dimensional fluctuations at $\delta=0$ to significant two-dimensional critical fluctuations can occur over a broad temperature range for $\delta\geq 0.1$ [44]. For each value of the phononic term $\delta$ that can be related to the term $\epsilon$, we can reliably determine the fluctuation term across a broad temperature range.
Accordingly, more specifically in Eq. (21), the four-dimensional quadratic coefficient highlights another temperature dependence, $r_{4}(T)\propto(T-T_{c})$, with a coefficient of proportionality that is different from $r_{0}$ and is also temperature-dependent. Consider a homogeneous superconductor with no superconducting current, and taking into account the minimization condition of the free energy in Eq. (14), the critical behavior of susceptibility length and order parameter below $T_{c}$ is characterized as follows:
$\chi\sim\Big|{\frac{T-T_{c}}{1-\epsilon\frac{T}{T_{c}}}}\Big|^{-1/2},$ (36)
and
$\phi\sim\Big|{\frac{T_{c}-T}{1-\epsilon\frac{T}{T_{c}}}}\Big|^{1/2}.$ (37)
Of central importance, it is noteworthy that in the function $v(T)=\Big|{\frac{T-T_{c}}{1-\epsilon\frac{T}{T_{c}}}}\Big|$, the parameter $\epsilon$ plays a primordial role depending on whether its value is equal to, greater than or less than 1. For $\epsilon=1$, we obtain $r_{4D}(T)=r_{0}$ below the critical point. In such a case, all thermodynamical quantities are constants, and the notion of fluctuation loses all its meaning in accordance with Landau’s theory for $d\geq 4$. For $\epsilon>1$, here again, we find the fundamental principles of the Ginzburg-Landau theory with an OP that increases continuously from zero starting at $T_{c}$. On the other hand, a whole new physics appears when $\epsilon<1$, thermodynamical functions as shown in Fig. 5 and Fig. 6 can both cancel out and diverge depending on the temperature below the critical point. In physics, particularly in phase transitions that involve abrupt and discontinuous changes in the properties of systems, some functions can take the value zero at certain points and diverge, tending to infinity at other specific points. This can occur in various contexts, such as electric potentials, gravitational fields, or oscillatory phenomena. In thermodynamics, the Gibbs free energy can exhibit both cancellation and divergence at certain points that often correspond to phase transitions or critical points in the system’s behavior. The cancellation of a wave function at the boundaries of a potential well is a boundary condition that ensures that the particle is confined within the well. The divergence outside the well is a mathematical consequence of the infinite nature of the potential. The cancellation and the divergence of functions in physics are often related to the nature of the forces or fields involved or to how these phenomena are modeled.
Figure 5: Illustration of the susceptibility as a function of temperature in a four-dimensional system, considering various values of the scaled fluctuation-correction term, $\epsilon$. For a critical temperature of $T_{c}$ = 2.5 K, the condition $\epsilon<1$ indicates the relative distance between the cancellation point and the divergence point.
Figure 6: Illustration of the order parameter as a function of temperature in a four-dimensional system, considering various values of the scaled fluctuation-correction term, $\epsilon$. For a critical temperature of $T_{c}$ = 2.5 K, the condition $\epsilon<1$ indicates the relative distance between the cancellation point and the divergence point.
Our model specifically applies to systems that have multiple competing order parameters. In certain physical systems, opposing order parameters such as electric and magnetic orders can partially cancel each other, leading to a complex overall ordering. For example, in $SU(N)$ gauge theory, a complex order parameter can be broken down into real and imaginary components. When $N\geq 3$, the various components may partially cancel each other, resulting in a complex form of ordering rather than a simpler one. A physical system that demonstrates both divergence and cancellation of an order parameter is often observed in high-energy physics. In these cases, divergences are addressed through a process known as renormalization. Similarly, in complex systems near a quantum critical point, competing order parameters can lead to cancellations. For example, in some theoretical models, divergences from one loop diagram may be canceled by divergences from other diagrams, resulting in a finite outcome.
This specific case with $\epsilon<1$ will be examined in greater detail elsewhere, in the context of a gravitational phase transition where the significance of dimension four is fully realized. Slow first-order phase transitions [38] can generate significant inhomogeneity, potentially leading to the formation of primordial black holes. In scenarios involving strong gravitational fields or high densities, such as in the early universe or within compact astrophysical objects, gravitational forces can significantly influence phase transitions. These influences can arise from pressure variations due to gravity, alterations in the effective potential of a system, or even through the generation of gravitational waves during the phase transition
IV Results and discussions
$\Delta{C_{0}}(d,T)=\frac{r^{2}_{0}T}{4b_{0}}\Bigg[\bigg(1+T_{c}\frac{\partial{\kappa(d,T)}}{\partial{T}}\bigg)^{2}+T^{2}_{c}\Big(\tau+\kappa(d,T)\Big)\frac{\partial^{2}{\kappa(d,T)}}{{\partial{T}}^{2}}\Bigg].$ (38)
In the expression of Eq. (38), when $\Big(\tau+\kappa(d,T)\Big)\Big|_{T=T_{c_{d}}}$ = 0 vanishes at the renormalized transition point, the quantity $\bigg(1+T_{c}\frac{\partial{\kappa(d,T)}}{\partial{T}}\Big|_{T=T_{c_{d}}}\bigg)^{2}$ provides information about the magnitude of the jump in the specific heat at that same point. If there are significant contributions from scattering states, the height of this jump can either increase or decrease substantially. Additionally, the interaction between $r_{0}(T-T_{c})$ and $\Omega(d,T)$ may result in inhomogeneities in the sample, which can potentially lead to the disappearance of the specific heat jump altogether. Hence, the anomalous part of the specific heat at the renormalized critical point is defined by this relationship.
By setting $\epsilon=\frac{\eta}{r_{0}T_{c}}$, we can easily establish the following result
$\Delta{C_{0}}(d,T_{c_{d}})=\left\{\begin{array}[]{llll}\vskip 8.5359pt0,&\textrm{for}\phantom{...}d=1\\ \vskip 8.5359pt(1+\epsilon)\Delta{C_{0}},&\textrm{for}\phantom{...}d=2\\ \vskip 8.5359pt0\phantom{..}\textrm{or}\phantom{..}4(1+\epsilon^{2})^{2}\Delta{C_{0}},&\textrm{for}\phantom{...}d=3\\ \vskip 8.5359pt(1-\epsilon)^{-2}\Delta{C_{0}},&\textrm{for}\phantom{...}d=4.\end{array}\right$. (44)
where $\Delta{C_{0}}=\frac{r^{2}_{0}T_{c}}{4b_{0}}$ is the mean-field specific heat jump. The renormalized specific heat jump is influenced by the system’s dimension and intrinsic parameters. Self-consistent calculations of Eq. (16) yield corrections that describe either a disappearance, linear growth, or quasi-exponential growth of the specific heat jump at the renormalized critical point. Linear growth is observed in two-dimensional systems, while exponential growth of the specific heat jump is seen in three-dimensional systems. Notably, one-dimensional systems do not exhibit a specific heat jump, and three-dimensional systems may also lack this jump under certain conditions. In fact, by solving Eq. (16) for $d$ = 3, we find two solutions; one of these leads to the absence of the specific heat jump at the critical temperature. This phenomenon occurs due to the competition between $r_{0}(T-T_{c})$ and $\Omega(d,T)$, which results in competing phases within the system and, consequently, a competition between the resulting order parameters. Additionally, the disappearance of the specific heat jump in superconducting systems can be attributed to a crossover in the type of order parameters, as discussed in Refs. [14], which leads to a significant loss of entropy at the critical temperature [13, 32]. Exponential growth of the specific heat jump has been observed in families of $\beta$-pyrochlore oxide superconductors [33, 34, 35], which can be explained by a significant lattice contribution coupled with strong fluctuations in the order parameter. However, certain families of high-$T_{c}$ superconductors, such as TlBaCaCuO and BiSrCaCuO, do not exhibit a conventional specific heat jump at the critical temperature ($T_{c}$) [15]. This lack of a clear jump can be attributed to the presence of a partial gap in the electronic density of states at the Fermi level [36]. Another potential explanation involves the smearing of $T_{c}$ due to sample inhomogeneities and the competition or interplay between superconductivity and other electronic phases, such as antiferromagnetism. Qualitatively, the results from Eq. (44) suggest that a small value of $\eta$ can either cancel out or significantly enhance the jump in specific heat at the critical temperature.
V Conclusion
Acknowledgments
Appendix A Hartree-Fock derivation and treatment of the correction-term $\Omega(d,T)$
To provide a comprehensive understanding, we present the main results related to the self-consistent correction term, denoted as $\Omega(d,T)$. We discussed these approximations in our previous work [25], but we did not specify their dimensional expressions. These approximations have been applied in various physical systems [28, 27]. Indeed, in the continuum limit, the Ginzburg-Landau theory provides a generalized partition function $Z$, which is derived by integrating over all possible configurations $\mathcal{D}\phi(\textbf{r})$ of the order parameter [31].
@@ Appendix A Hartree-Fock derivation and treatment of the correction-term $\Omega(d,T)$
More importantly, as important point in our treatment the correlation function relating the order parameter and density at different positions is given by
@@
In this context, for more information on calculating thermodynamic functions, we refer readers to Ref. [7]. Accordingly, one can introduce the Fourier modes through the decomposition of the spatially dependent order parameter $\phi(\textbf{r})$ into its Fourier component $\phi(\textbf{q})$ and vice-versa, and which are defined as follows:
$\phi(\textbf{q})=\frac{1}{V}\int d^{d}\textbf{r}\phi(\textbf{r})\exp\{-i\textbf{q}\cdot\textbf{r}\},$
$\phi(\textbf{r})=\frac{V}{(2\pi)^{d}}\int d^{d}\textbf{q}\phi(\textbf{q})\exp\{i\textbf{q}\cdot\textbf{r}\}.$ (49)
Using the Fourier components and within the framework of the theoretical Landau mean-field theory in the absence of an external field, the corresponding partition function (normalization) obtained by integration over all possible configurations of the order parameter is obtained as:
$\mathcal{Z}=\int\mathcal{D}\phi\exp\left[-\beta\left\{\sum_{q}\big(r_{0}(T-T_{c})+\mathcal{C}q^{2}\big)\phi(\textbf{q})\phi(-\textbf{q})+\frac{b_{0}}{L^{d}}\sum_{{q_{i}}}\phi(\textbf{q}_{1})\phi(\textbf{q}_{2})\phi{\textbf{q}_{3}}\phi(-\textbf{q}_{1}-\textbf{q}_{2}-\textbf{q}_{3})\right\}\right]$ (50)
where $\mathcal{D}\phi(\textbf{q})$ is an integration measure for real modes representing all possible configurations of the order parameter given by
$\mathcal{D}\phi=\prod_{q}(2\pi)^{-1}d\phi(\textbf{q})d\phi(-\textbf{q}).$ (51)
In this respect, $\int\mathcal{D}\phi(\textbf{q})$ defines the functional integral of real modes. To describe a finite-temperature phase transition, we consider a set of weakly interacting particles or quai-particles. By associating the $\phi^{2}$ term with essential fluctuations and linking the $\phi^{4}$ term to redundant fluctuations, the quartic term acts as an interaction term among the Fourier components of the order parameter. These components represent the fluctuation modes of the order parameter and can be decoupled into a product of two quantities as
$\sum_{q_{i}}\phi(\textbf{q}_{1})\phi(\textbf{q}_{2})\phi(\textbf{q}_{3})\phi(-\textbf{q}_{1}-\textbf{q}_{2}-\textbf{q}_{3})\approx 6\sum_{q}\left(\sum_{q^{\prime}<\Lambda}\langle|\phi(\textbf{q}^{\prime})\phi(-\textbf{q}^{\prime})|\rangle\right)\phi(\textbf{q})\phi(-\textbf{q}).$ (52)
Thus, the approximation $|\phi(q)|^{4}\approx 6\langle|\phi(q)|^{2}\rangle|\phi(q)|^{2}$ is used, which assumes that the Fourier components interact only through the mean field produced by other modes. The factor of 6 here accounts for all possible contractions and $\Lambda$ is a cutoff value corresponding to a lattice periodicity. According to the Hartree-Fock [39, 40] approximation, low-dimensional systems are unable to establish long-range order at finite temperatures because of fluctuations in the order parameter. Consequently, fluctuations with wavelengths greater than the coherence length of the sample are considered negligible. This decoupling is justified when the quartic coefficient of the expansion is small.
By making the following replacement:
$|\phi(q)|^{4}\rightarrow 6\langle\phi(q)^{2}\rangle|\phi(q)|^{2},$ (53)
in the partition function, the functional integral can be evaluated analytically when dealing with a Gaussian field. By considering the expected fact that the expectation value $\langle\phi(q)^{2}\rangle$ does not explicitly depend on the wave number $q$, we can subsequently derive according to Eqs. (A1), (A2), (A3), (A4) and (A6) the self-consistent equation for $\Omega(d,T)$
$\Omega(d,T)=\frac{3b_{0}}{V}\langle\phi(q)^{2}\rangle=\frac{3b_{0}}{V}\left[2\beta\sum_{q}\left(\mathcal{C}q^{2}+r_{0}(T-T_{c})+\Omega(d,T)\right)\right]^{-1}.$ (54)
By making a transition to the continuum through the relation
$\sum_{q}(...)=L^{d}(2\pi)^{-d}\int_{-\infty}^{+\infty}(...)d^{d}q,$ (55)
we can therefore derive the self-consistent KMDDH equation of the quadratic correction coefficient $\Omega(d,T)$ given as
$\Omega(d,T)-\eta\left(\frac{T}{T_{c}}\right)\left(\tau+\frac{\Omega(d,T)}{r_{0}T_{c}}\right)^{\frac{d}{2}-1}=0$ (56)
where $\tau=(T-T_{c})/T_{c}$ and $\eta$ denotes the dimension-dependent variational energetic parameter related to the harmonic variance of the order parameter. One can therefore make the Gaussian approximation with the following transformation for the quadratic coefficient of the Ginzburg-Landau free energy functional as follows:
$r_{0}(T-T_{c})\rightarrow r_{0}(T-T_{c})+\Omega(d,T)$ (57)
which now takes into account fluctuations and correlations of the order parameter. These results from the Hartree-Fock approximation show both qualitative and quantitative agreement with specific experimental facts, as described by the Mermin-Wagner-Hohenberg theorem. These results, represented by dashed curves in Fig. 3, pertain specifically to one-dimensional systems and two-dimensional system for large value of $\eta$. The parameter $\eta$ is dimension-dependent and accounts for scattering states contributions [25].